OpenStudy (anonymous):

Find limit as x -> infinite, ((4 + (9x+1)^2)/(500x + 3x^2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{4+81x^2+18x+1}{500x+3x^2}\] \[\lim_{x \rightarrow \infty} \frac{81x^2+18x+5}{3x^2+500x}=\frac{81}{3}=27\]

OpenStudy (campbell_st):

expand and simplify the numerator then \[\lim_{x \rightarrow \infty}(81x^2 + 28x +5)/(500x + 3x^2)\] then divide every term by x^2 will result in 81/3 as all other terms tend to zero as x approaches infinity so the limit will be 27